# JWST Observations of the Double Nucleus in NGC 4486B: Possible Evidence for a Recent Binary SMBH Merger and Recoil

### v1

@@ 3.2 Constraints on the Recoil Kick and SMBH Binary Parameters
Figure 4: Expected GW recoil kick magnitude, $V_{\rm{kick}}$, as a function of the pre-merger black hole mass ratio, $q$. The thick purple line shows the mean, and the shaded region a $1\sigma$ confidence interval estimated via Monte Carlo simulations. The recoil kick required to induce the double nucleus of NGC 4486B is shown in the gray horizontal dashed line, and the mass of the corresponding pre-merger companion to NGC 4486B’s central SMBH is indicated on the top horizontal axis. The gray vertical dashed line marks $q=0.15$, the lower-limit pre-merger black hole mass ratio for NGC 4486B. Adapted from Akiba.2025 which uses the analytical prescription from Lousto.2010; Lousto.2012.
@@
Akiba.2021 simulated disks with stars initially in circular orbits about an SMBH. Following an in-plane GW recoil kick imparted on the black hole, the stars preferentially align their eccentricity vectors perpendicular to the recoil kick with magnitude
@@
where $v_{\rm{kick}}$ is the recoil kick speed and $v_{c}$ is the initial average circular speed of the stars. In their simulations, they found that immediately after the kick, the mean eccentricity jumps to $|\vec{e}_{\rm{avg}}|\sim 1$ but subsequently decreases via large amplitude oscillations to smaller values. Simulations show that while the disk gradually diffuses in semi-major axis and orientation (as $|\vec{e}_{\rm{avg}}|$ decreases), $R_{c}$, the characteristic radius of the disk and the degree of apsidal alignment remain stable over $\gtrsim 10$ Myr (Akiba.2021; Akiba.2023). Akiba.2023 noticed that by setting $|\vec{e}_{\rm{avg}}|\,\sim 1$, one can determine $R_{c}$ of the resulting eccentric disk as,
@@
In Figure 4, we show the expected recoil kick magnitude as a function of the pre-merger black hole mass ratio, $q=m_{2}/M_{1}$ (where $M_{1}$ is the more massive black hole in the binary); adapted from Akiba.2025 applying the analytical prescription given in Lousto.2010; Lousto.2012; Fragione.2023. The light purple region shows the Monte Carlo results of $10^{6}$ evaluations of the analytical model (described in Fragione.2023) with randomized spins. The dimensionless spins for both black holes are drawn from an isotropic distribution in direction and a uniform distribution in magnitude ([0, 1]). The thick purple line shows the average kick magnitude as a function of $q$ marginalized over the spin distribution. The mean and $1\sigma$-confidence interval of the recoil kick distribution is shown, and the estimated $V_{\mathrm{kick}}\sim 340\,\mathrm{km\ s}^{-1}$ , required to induce NGC 4486B’s double nucleus is indicated with the dashed gray line. This recoil kick implies a pre-merger mass ratio of $q\approx 0.24$ which corresponds to a predicted companion mass of $7\times 10^{7}\,M_{\odot}$ (the calculation does not account for mass lost to GWs). However, the estimated recoil kick magnitude falls within one standard deviation of the mean for all mass ratios $q>0.15$, so one can only determine a lower limit at the companion mass of $4.7\times 10^{7}\,M_{\odot}$. Additionally, it is important to note that these recoil kick and mass ratio estimates assume that stars are initially on circular orbits. Akiba.2023 showed that the inclusion of initial eccentricities reduces the kick magnitude required to produce an eccentric disk at a given characteristic radius, so a degeneracy exists between the companion mass estimate and the pre-merger stellar eccentricities in the nuclear star cluster. This analytic framework establishes the connection between the eccentric disk geometry and the GW recoil kick. We emphasize that the purpose of this analytic treatment is to assess the plausibility of this scenario. More careful calculations and simulations will be carried out in the future. We build upon the above analysis in the following sections to explore the plausible timescales of such an event and the dynamical signatures it may have imprinted on the double nucleus in NGC 4486B.
@@
4 Post kick trajectory
@@
Considering the strong evidence in support of a recoil kick, we now explore the possible post-kick trajectories of the SMBH. We construct an $N$-body representation of the best-fit Schwarzchild model of NGC 4486B with $10^{6}$ particles. This model, which has no dark matter halo, has a best-fit black hole mass of $M_{\rm BH}=3.51\times 10^{8}\mbox{$M_{\odot}$}$, total stellar mass $M_{*}=7.44\times 10^{9}\mbox{$M_{\odot}$}$, and mass-to-light ratio $\Upsilon_{[F850LP]}=4.1$. We introduce a black hole particle at the center of the system and give it a kick velocity $V_{\rm kick}$ along several directions. We explore $V_{\rm kick}=0.2V_{\rm esc}=154\,\mathrm{km\ s}^{-1}$, $V_{\rm kick}=0.5V_{\rm esc}=384\,\mathrm{km\ s}^{-1}$, and $V_{\rm kick}=0.75v_{\rm esc}=576\,\mathrm{km\ s}^{-1}$, where $V_{\rm esc}$ is the escape velocity from the galaxy center (the middle value being close to the kick velocity estimated in Section 3.2).
@@
The simulations are run using the publicly available code GyrfalcOn (Dehnen2002). Each particle is softened using the P1 softening kernel [^1] A modification of the Plummer softening kernel whose density is $\propto(1+x^{2})^{-7/2}$, with softening length 0.79 pc for the stars and 0.395 pc for the BH. We use hierarchical timesteps, where the timestep of each particle depends on it softening length $\epsilon$ and instantaneous acceleration $a_{i}$ as $\Delta t\sim 0.025\sqrt{\epsilon/\left|a_{i}\right|}$. According to this criterion, a circular orbit around the black hole at 0.79 pc (the softening length) is resolved with at least $\sim$100 timesteps.
@@
Figure 5 shows properties of the resulting trajectory of the SMBH in the various cases. The top and bottom panels show the SMBH’s distance and velocity with respect to the galaxy’s center of mass [^2] In order to calculate the center of mass, we use the $90\%$ most bound stars in the simulation, as some stars may get unbound during the recoil kick. The results differ only slightly based on the exact center of mass definition., and the right, middle, and left panels correspond to cases with $V_{\rm kick}=0.2V_{\rm esc}$, $0.5V_{\rm esc}$, and $0.75V_{\rm esc}$, respectively. Notably, in all simulations, the line-of-sight velocity dispersion peak consistently follows the SMBH, justifying our assumption in Section 3.1 that the off-center velocity dispersion peak marks the location of the SMBH. Within each panel, the various curves correspond to different initial kick directions. In all panels, the dotted black line marks the observed positional offset of the SMBH (6 pc) and the localized $v_{\mathrm{los}}$ offset (16 km s-1) measured at the location of the SMBH (see top panel of Fig. 2). This feature is assumed to trace the mean motions of stars bound to the SMBH but may not represent the intrinsic 3D velocity of the black hole itself. However, for comparison with the simulations, we refer to this quantity as the “observed line-of-sight velocity”. The black dashed lines in the lower and upper panels show the expected velocity and position amplitudes due to Brownian motion, respectively, which are given by Gualandris.2008:
@@
In every trajectory shown in Figure 5, the black hole rapidly sinks back toward the galaxy center, reaching amplitudes comparable to the observed positional and $v_{\mathrm{los}}$ offsets, after which these quantities fall below the observed values. The return timescale increases with kick strength, from $\sim$10 Myr for $0.2V_{\rm esc}$ to $\sim$30 Myr for $0.5V_{\rm esc}$ and $\sim$80 Myr for $0.75V_{\rm esc}$. Although these timescales differ, all are extremely short relative to a Hubble time. Furthermore, the observed $v_{\mathrm{los}}$ offset is only one component of the possible 3D velocity of the SMBH. At late times, the black hole retains a residual velocity relative to the galaxy center that is clearly resolved above the expected Brownian amplitude. This residual motion is likely the result of a core-stalling effect (e.g., Read2006), but it remains insufficient to account for the observed velocity offset. Interestingly, the timescale for the black hole to return to the center is longer if the kick is along the Z-axis, compared to the other directions, although this timescale is still significantly shorter than a Hubble time. This will be further investigated in a follow-up paper (Jha et al, in preparation).
@@
5 Alternative Dynamical Scenarios for Black Hole Displacement
5.1 Buoyancy-Induced Black Hole Offset
Dynamical buoyancy is the opposite of dynamical friction, where the host system exerts a net positive (outward) torque on a massive perturber, leading to its outward motion (Cole2012; Banik2021; Banik2022; Dattathri2025b). This phenomenon is typically associated with cored density profiles. Since NGC 4486B has a central core (Figure 1), we investigate here whether dynamical buoyancy operates, and if so, what the associated offsets are.
We initialize an $N$-body simulation with the same setup and parameters as Section 4, but without any recoil kick, i.e., the SMBH is initially at rest at the galaxy center. We then let the system evolve and track the SMBH’s motion around the galaxy center. Figure 6 shows the resulting trajectory of the SMBH’s position (top) and velocity (bottom). Similar to Figure 5, the black dashed line shows the expected Brownian motion amplitude, and the black dotted line shows the observed offsets. The SMBH shows a clear initial outward motion even without a recoil kick. This motion is well resolved above the Brownian amplitude, indicating that it is not a numerical artifact. The offset reaches a maximum amplitude of $\sim 50\,\mathrm{km\ s}^{-1}$ and $4$ pc away from the center at $T\sim 2$ Myr. However, similar to the recoil simulations (Figure 5), the black hole quickly sinks back to the center. After $T\gtrsim 15$ Myr, the SMBH is orbiting around the galaxy center at a speed of $\sim 5$$\mathrm{km\ s}^{-1}$and a distance 121-2 pc, which is significantly lower than the observed offsets.
The origin of the buoyancy in our system is unclear. Dattathri2025b suggested that in spherical isotropic systems, dynamical buoyancy is due to an unstable dipole mode in the system. This instability may arise if the distribution function has a region where ${\rm d}f/{\rm d}E>0$ (Dattathri2025a). However, NGC 4486B is not isotropic and has a net rotation (Figure 2); hence, it is unclear whether the same instability would occur here. A recent study by Boily2025 analyzed the motion of BHs in anisotropic galaxy cores and found that if the system has an excess of streaming motion over its velocity dispersion, a central black hole can gain angular momentum and move outward. Both of the above processes could work in tandem to produce an offset SMBH. However, regardless of the mechanism, it is clear from Figure 6 that if the SMBH in NGC 4486B was initially at rest in the center, it could attain a significant velocity offset due to dynamical buoyancy for a short period of time $\sim 6-7$Myr, but the position offset is always well below the observed value of 6 pc. Note that the observed 6 pc separation is a projected distance; the true 3D offset can only be larger, which further strengthens this argument.
@@ 5.1 Buoyancy-Induced Black Hole Offset
5.2 Pre-Merger Binary Black Hole in Prograde Orbit
@@
Although NGC 4486B is overall a dispersion-dominated galaxy, it exhibits a considerable central rotation ($|v|/\sigma\simeq$ 0.40-0.65). Studies have shown that even this level of rotational support can profoundly influence the dynamics of SMBH binaries in rotating hosts (holley+15; Rasskazov+16; Mirza+17; Khan2020). In particular, simulations demonstrate that when a binary is on a prograde orbit aligned with the stellar rotation, its center-of-mass (CoM) can settle into a stable orbit around the galactic center with a characteristic radius comparable to the influence radius of the two BHs. Once established, such a system can survive for extended periods, maintaining a corotating orbit while the SMBHs in binary continue to shrink their relative separation, ultimately coalescing following GW emission.
This scenario is especially relevant for NGC 4486B. Photometric and kinematic studies indicate that it is an overall old, dynamically relaxed galaxy: its stellar populations are predominantly ancient, its large-scale light distribution is smooth, and no evidence exists for recent large-scale tidal disturbances (e.g., Ferrarese.2006; Guerou.2015; Janz.2016). Thus, any major merger that contributed to its present structure likely occurred several gigayears ago (Khan2018; Kelly2025), consistent with its environment deep within the Virgo cluster. Nevertheless, if that merger brought in a secondary SMBH on a prograde trajectory, the resulting binary could have settled into a long-lived corotating orbit before the SMBHs achieved a coalescence. This outcome naturally extends the effective lifetime of the binary off-center phase and is supported by the $N$-body simulations of rotating systems cited above.
@@
Once again, our initial $N$-body model reflects the current structure of NGC 4486B rather than that of its progenitor, which likely possessed a steeper central density cusp prior to binary scouring. Even so, the simulation demonstrates that in a host galaxy with the modest but non-negligible large-scale rotation observed, a forming SMBH binary can become trapped in a corotation resonance for a finite period. Because of the high computational cost of direct $N$-body integrations, we have evolved this model only to $\sim 30$ Myr. Nevertheless, our results indicate that while NGC 4486B is globally an old and relaxed system, its central SMBH binary could plausibly have remained in a long-lived resonant, corotating configuration following a galaxy–galaxy merger in the cluster outskirts (the Virgo cluster crossing time is $\sim 10^{9}$ yr), and may have undergone coalescence only relatively recently. Based on our estimates, the SMBH binary would require $\approx 500~\mathrm{Myr}$ to coalesce after forming via dynamical-friction–driven inspiral—an inspiral that itself may take several hundred Myr if the secondary SMBH arrives at a separation of $\sim 1~\mathrm{kpc}$ during the preceding merger.

### v2

@@ III.2 Constraints on the Recoil Kick and SMBH Binary Parameters
Figure 4: Expected GW recoil kick magnitude, $V_{\rm{kick}}$, as a function of the pre-merger black hole mass ratio, $q$. The thick purple line shows the mean, and the shaded region a $1\sigma$ confidence interval estimated via Monte Carlo simulations. The recoil kick required to induce the double nucleus of NGC 4486B is shown in the gray horizontal dashed line, and the mass of the corresponding pre-merger companion to NGC 4486B’s central SMBH is indicated on the top horizontal axis. The gray vertical dashed line marks $q=0.15$, the lower-limit pre-merger black hole mass ratio for NGC 4486B. Adapted from Akiba et al. (2025) which uses the analytical prescription from Lousto et al. (2010, 2012).
@@
Akiba & Madigan (2021) simulated disks with stars initially in circular orbits about an SMBH. Following an in-plane GW recoil kick imparted on the black hole, the stars preferentially align their eccentricity vectors perpendicular to the recoil kick with magnitude
@@
where $V_{\rm{kick}}$ is the recoil kick speed and $v_{c}$ is the initial average circular speed of the stars. In their simulations, they found that immediately after the kick, the mean eccentricity jumps to $|\vec{e}_{\rm{avg}}|\sim 1$ but subsequently decreases via large amplitude oscillations to smaller values. Simulations show that while the disk gradually diffuses in semi-major axis and orientation (as $|\vec{e}_{\rm{avg}}|$ decreases), $R_{c}$, the characteristic radius of the disk and the degree of apsidal alignment remain stable over $\gtrsim 10$ Myr (Akiba & Madigan, 2021, 2023). Akiba & Madigan (2023) noticed that by setting $|\vec{e}_{\rm{avg}}|\,\sim 1$, one can determine $R_{c}$ of the resulting eccentric disk as,
@@
In Figure 4, we show the expected recoil kick magnitude as a function of the pre-merger black hole mass ratio, $q=m_{2}/M_{1}$ (where $M_{1}$ is the more massive black hole in the binary). The kick magnitudes are computed using analytic fitting formulae calibrated to numerical relativity simulations of binary black hole mergers (Lousto et al., 2010, 2012), as implemented in the Monte Carlo framework of Fragione & Rasio (2023) and applied to END scenarios by Akiba et al. (2025). The light purple region shows the Monte Carlo results of $10^{6}$ evaluations of the analytical model (described in Fragione & Rasio (2023)) with randomized spins. The dimensionless spins for both black holes are drawn from an isotropic distribution in direction and a uniform distribution in magnitude ([0, 1]). The thick purple line shows the average kick magnitude as a function of $q$ marginalized over the spin distribution. The mean and $1\sigma$-confidence interval of the recoil kick distribution is shown, and the estimated $V_{\mathrm{kick}}\sim 340\,\mathrm{km\ s}^{-1}$ , required to induce NGC 4486B’s double nucleus is indicated with the dashed gray line. This recoil kick implies a pre-merger mass ratio of $q\approx 0.24$ which corresponds to a predicted companion mass of $7\times 10^{7}\,M_{\odot}$ (the calculation does not account for mass lost to GWs). However, the estimated recoil kick magnitude falls within one standard deviation of the mean for all mass ratios $q>0.15$, so one can only determine a lower limit at the companion mass of $4.7\times 10^{7}\,M_{\odot}$. Additionally, it is important to note that these recoil kick and mass ratio estimates assume that stars are initially on circular orbits. Akiba & Madigan (2023) showed that the inclusion of initial eccentricities reduces the kick magnitude required to produce an eccentric disk at a given characteristic radius, so a degeneracy exists between the companion mass estimate and the pre-merger stellar eccentricities in the nuclear star cluster. This analytic framework establishes the connection between the eccentric disk geometry and the GW recoil kick.
We emphasize that this analytic treatment is intended to provide order-of-magnitude constraints and a consistency check on the recoil scenario, rather than a fully predictive, galaxy-specific model. A direct comparison between theory and the detailed observables of NGC 4486B—such as generating surface-brightness and two-dimensional kinematic maps from tailored $N$-body simulations—is beyond the scope of the present paper. Existing END simulations are intentionally idealized, typically modeling isolated eccentric disks with limited particle numbers to capture the underlying physical mechanism and viewing-angle dependencies, whereas a realistic forward model for NGC 4486B would require high-resolution simulations that self-consistently include both the stellar bulge and the nuclear disk and explore a broad parameter space (disk structure, eccentricity distribution, kick amplitude and direction, and viewing angle). We therefore focus here on establishing plausibility and placing data-driven constraints on the recoil kick scale using analytic scalings and the observed kinematic signatures, and view galaxy-specific modeling as a natural next step to be pursued in future work.
We build upon the above analysis in the following sections to explore the plausible timescales of such an event and the dynamical signatures it may have imprinted on the double nucleus in NGC 4486B.
@@
IV Post-Kick Trajectory of the Recoiling SMBH
@@
Considering the strong evidence in support of a recoil kick, we now explore the possible post-kick trajectories of the SMBH. We construct an $N$-body representation of the best-fit Schwarzschild model of NGC 4486B with $10^{6}$ particles. This model, which has no dark matter halo, has a best-fit black hole mass of $M_{\rm BH}=3.51\times 10^{8}\mbox{$M_{\odot}$}$, total stellar mass $M_{*}=7.44\times 10^{9}\mbox{$M_{\odot}$}$, and mass-to-light ratio $\Upsilon_{[F850LP]}=4.1$. We introduce a black hole particle at the center of the system and give it a kick velocity $V_{\rm kick}$ along several directions. We explore $V_{\rm kick}=0.2V_{\rm esc}=154\,\mathrm{km\ s}^{-1}$, $V_{\rm kick}=0.5V_{\rm esc}=384\,\mathrm{km\ s}^{-1}$, and $V_{\rm kick}=0.75V_{\rm esc}=576\,\mathrm{km\ s}^{-1}$, where $V_{\rm esc}$ is the escape velocity from the galaxy center (the middle value being close to the kick velocity estimated in Section III.2).
@@
The simulations are run using the publicly available code GyrfalcOn (Dehnen, 2002). Each particle is softened using the P1 softening kernel [^1] A modification of the Plummer softening kernel whose density is $\propto(1+x^{2})^{-7/2}$, with softening length 0.79 pc for the stars and 0.395 pc for the BH. We use hierarchical timesteps, where the timestep of each particle depends on its softening length $\epsilon$ and instantaneous acceleration $a_{i}$ as $\Delta t\sim 0.025\sqrt{\epsilon/\left|a_{i}\right|}$. According to this criterion, a circular orbit around the black hole at 0.79 pc (the softening length) is resolved with at least $\sim$100 timesteps.
@@
Figure 5 shows properties of the resulting trajectory of the SMBH in the various cases. The top and bottom panels show the SMBH’s distance and velocity with respect to the galaxy’s center of mass [^2] In order to calculate the center of mass, we use the $90\%$ most bound stars in the simulation, as some stars may get unbound during the recoil kick. The results differ only slightly based on the exact center of mass definition., and the right, middle, and left panels correspond to cases with $V_{\rm kick}=0.2V_{\rm esc}$, $0.5V_{\rm esc}$, and $0.75V_{\rm esc}$, respectively. Notably, in all simulations, the line-of-sight velocity dispersion peak consistently follows the SMBH, justifying our assumption in Section III.1 that the off-center velocity dispersion peak marks the location of the SMBH. Within each panel, the various curves correspond to different initial kick directions. In all panels, the dotted black line marks the observed positional offset of the SMBH (6 pc) and the localized $v_{\mathrm{los}}$ offset (16 km s-1) measured at the location of the SMBH (see top panel of Fig. 2). This feature is assumed to trace the mean motions of stars bound to the SMBH but may not represent the intrinsic 3D velocity of the black hole itself. However, for comparison with the simulations, we refer to this quantity as the “observed line-of-sight velocity”. The black dashed lines in the lower and upper panels show the expected velocity and position amplitudes due to Brownian motion, respectively, which are given by Gualandris & Merritt (2008):
@@
In every trajectory shown in Figure 5, the black hole rapidly sinks back toward the galaxy center, reaching amplitudes comparable to the observed positional and $v_{\mathrm{los}}$ offsets, after which these quantities fall below the observed values. The return timescale increases with kick strength, from $\sim$10 Myr for $0.2V_{\rm esc}$ to $\sim$30 Myr for $0.5V_{\rm esc}$ and $\sim$80 Myr for $0.75V_{\rm esc}$. Although these timescales differ, all are extremely short relative to a Hubble time. Furthermore, the observed $v_{\mathrm{los}}$ offset is only one component of the possible 3D velocity of the SMBH. At late times, the black hole retains a residual velocity relative to the galaxy center that is clearly resolved above the expected Brownian amplitude. This residual motion is likely the result of a core-stalling effect (e.g., Read et al., 2006), but it remains insufficient to account for the observed velocity offset. Interestingly, the timescale for the black hole to return to the center is longer if the kick is along the Z-axis, compared to the other directions, although this timescale is still significantly shorter than a Hubble time. This rapid damping of the SMBH motion is consistent with recent work showing that the long-term trajectories of massive perturbers in galaxy cores depend sensitively on the core structure and are generally damped on relatively short timescales in realistic potentials (Rawlings et al., 2025). This will be further investigated in a follow-up paper (Jha et al, in preparation).
@@
V Alternative Dynamical Scenarios for Black Hole Displacement
V.1 Buoyancy-Induced Black Hole Offset
Dynamical buoyancy is the opposite of dynamical friction, where the host system exerts a net positive (outward) torque on a massive perturber, leading to its outward motion (Cole et al., 2012; Banik & van den Bosch, 2021, 2022; Dattathri et al., 2025a). This phenomenon is typically associated with cored density profiles. Since NGC 4486B has a central core (Figure 1), we investigate here whether dynamical buoyancy operates, and if so, what the associated offsets are.
We initialize an $N$-body simulation with the same setup and parameters as Section IV, but without any recoil kick, i.e., the SMBH is initially at rest at the galaxy center. We then let the system evolve and track the SMBH’s motion around the galaxy center. Figure 6 shows the resulting trajectory of the SMBH’s position (top) and velocity (bottom). Similar to Figure 5, the black dashed line shows the expected Brownian motion amplitude, and the black dotted line shows the observed offsets. The SMBH shows a clear initial outward motion even without a recoil kick. This motion is well resolved above the Brownian amplitude, indicating that it is not a numerical artifact. The offset reaches a maximum amplitude of $\sim 50\,\mathrm{km\ s}^{-1}$ and $4$ pc away from the center at $T\sim 2$ Myr. However, similar to the recoil simulations (Figure 5), the black hole quickly sinks back to the center. After $T\gtrsim 15$ Myr, the SMBH is orbiting around the galaxy center at a speed of $\sim 5$$\mathrm{km\ s}^{-1}$and a distance 121-2 pc, which is significantly lower than the observed offsets.
The origin of the buoyancy in our system is unclear. Dattathri et al. (2025a) suggested that in spherical isotropic systems, dynamical buoyancy is due to an unstable dipole mode in the system. This instability may arise if the distribution function has a region where ${\rm d}f/{\rm d}E>0$ (Dattathri et al., 2025b). However, NGC 4486B is not isotropic and has a net rotation (Figure 2); hence, it is unclear whether the same instability would occur here. A recent study by Boily et al. (2025) analyzed the motion of BHs in anisotropic galaxy cores and found that if the system has an excess of streaming motion over its velocity dispersion, a central black hole can gain angular momentum and move outward. Both of the above processes could work in tandem to produce an offset SMBH. However, regardless of the mechanism, it is clear from Figure 6 that if the SMBH in NGC 4486B was initially at rest in the center, it could attain a significant velocity offset due to dynamical buoyancy for a short period of time $\sim 6-7$Myr, but the position offset is always well below the observed value of 6 pc. Note that the observed 6 pc separation is a projected distance; the true 3D offset can only be larger, which further strengthens this argument.
@@ V.1 Buoyancy-Induced Black Hole Offset
V.2 Pre-Merger Binary Black Hole in Prograde Orbit
@@
Although NGC 4486B is overall a dispersion-dominated galaxy, it exhibits a considerable central rotation ($|v|/\sigma\simeq$ 0.40-0.65). Studies have shown that even this level of rotational support can profoundly influence the dynamics of SMBH binaries in rotating hosts (Holley-Bockelmann & Khan, 2015; Rasskazov & Merritt, 2016; Mirza et al., 2017; Khan et al., 2020). In particular, simulations demonstrate that when a binary is on a prograde orbit aligned with the stellar rotation, its center-of-mass (CoM) can settle into a stable orbit around the galactic center with a characteristic radius comparable to the influence radius of the two BHs. Once established, such a system can survive for extended periods, maintaining a corotating orbit while the SMBHs in binary continue to shrink their relative separation, ultimately coalescing following GW emission.
This scenario is especially relevant for NGC 4486B. Photometric and kinematic studies indicate that it is an overall old, dynamically relaxed galaxy: its stellar populations are predominantly ancient, its large-scale light distribution is smooth, and no evidence exists for recent large-scale tidal disturbances (e.g., Ferrarese et al., 2006; Guérou et al., 2015; Janz et al., 2016). Thus, any major galaxy–galaxy merger that contributed to its present structure likely occurred several gigayears ago (Khan et al., 2018; Holley-Bockelmann et al., 2025), consistent with its environment deep within the Virgo cluster. Nevertheless, if that galaxy–galaxy merger brought in a secondary SMBH on a prograde trajectory, the resulting binary could have settled into a long-lived corotating orbit before the SMBHs achieved a coalescence. This outcome naturally extends the effective lifetime of the binary off-center phase and is supported by the $N$-body simulations of rotating systems cited above.
@@
Once again, our initial $N$-body model reflects the current structure of NGC 4486B rather than that of its progenitor, which likely possessed a steeper central density cusp prior to binary scouring. Even so, the simulation demonstrates that in a host galaxy with the modest but non-negligible large-scale rotation observed, a forming SMBH binary can become trapped in a corotation resonance for a finite period. Because of the high computational cost of direct $N$-body integrations, we have evolved this model only to $\sim 30$ Myr. Nevertheless, our results indicate that while NGC 4486B is globally an old and relaxed system, its central SMBH binary could plausibly have remained in a long-lived resonant, corotating configuration following a galaxy–galaxy merger in the cluster outskirts (the Virgo cluster crossing time is $\sim 10^{9}$ yr), and may have undergone coalescence only relatively recently. We estimated the supermassive black hole (SMBH) binary coalescence time using a semi-analytical prescription that accounts for orbital energy losses driven by both stellar dynamical interactions and GW emission (Khan et al., 2012; Sesana & Khan, 2015). The stellar hardening rate was directly measured from our simulations and incorporated into the model to capture the contribution from three-body encounters. In addition, we included the energy and angular momentum losses due to GW radiation following the formalism of Peters & Mathews (1963), who derived the GW energy loss for two point masses in orbit about one another. Based on our estimates, the SMBH binary would require $\approx 500~\mathrm{Myr}$ to coalesce after its formation and this pahase will be preceeded by dynamical-friction–driven inspiral that itself may take several hundred Myr if the secondary SMBH arrives at a separation of $\sim 1~\mathrm{kpc}$ during the preceding galaxy–galaxy merger.